OpenStudy (anonymous):

help me pls,,, marvin invested P6000 more at 5% than he invested at 8%. his total income was P2420 . how much did he invested at each rate?

OpenStudy (anonymous):

I think he invested 16308 at 8% and (16308+6000) at 5%. 16308 * 8% + (16308+6000) * 5% = 2420

OpenStudy (anonymous):

Might be a redundant question but should the amount of time be taken into consideration?

OpenStudy (anonymous):

what's the solution?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

the question does not make sense

OpenStudy (anonymous):

what is P1600 ?

OpenStudy (anonymous):

P stands for pounds?

OpenStudy (anonymous):

with a name like kyragayle, you must be a british

OpenStudy (anonymous):

no im not

OpenStudy (anonymous):

what the heck is P6000? is that dollars? P is what?

OpenStudy (anonymous):

marvin is a flutter up

OpenStudy (anonymous):

do we need that P ?

OpenStudy (anonymous):

i can answer this if we ignore the P , did you mean to write $ instead of P

OpenStudy (anonymous):

let x = amount invested at 8% y = amount invested at 5% it says that y = 6000 + x because the amount invested at 5% is 6000 more than the amount invested at 8%. so we have .05 ( 6000 + x ) + .08 x = 2420

OpenStudy (anonymous):

i get x = 16307.69

OpenStudy (anonymous):

i think she was using I=Prt, or Interest = Principle x rate x time, thus the total amount of interest = the original amount x rate of interest (usually in decimals) x the amount of time (months, years, etc). might be where the P came from.

OpenStudy (anonymous):

oh, then i answered a different question?

OpenStudy (anonymous):

nah i think you got it right, just sayin she was noting it was principle

OpenStudy (anonymous):

wait, does it change the answers? let me see

OpenStudy (anonymous):

wait, P6000 means P = 6000, i think

OpenStudy (anonymous):

the principal is 6000 , so you dont need your whole I = Prt thing

OpenStudy (anonymous):

because youre given P, r is also given, and t is probably over a year since its not given

OpenStudy (anonymous):

zhill, the time actually gets factored out

OpenStudy (anonymous):

i know, just pointing out the SHE may have been clarifying for the rest of the world.

OpenStudy (anonymous):

p is peso

OpenStudy (anonymous):

actually the time does not get factored out, this problem is ill defined

OpenStudy (radar):

Let x = amount of Peso invested at 8% then x+6000 = amount of Pesos invested at 5% .08x = amount of annual income from 8% investment .05(x+6000) = amount of annual income from 5% investment. .08x +.05(x+6000)=total annual income from total investments .08x+.05(x+6000)=2420 (Pesos) Solve this for x .08x + .05x + 300=2420 .13x=2420-300=2120 or .13x=2120 x=16307.69 peso at 8% 22307.69 pesos at 5%